Before a captive breeding program and subsequent reintroduction of the black-footed ferret it is believed that their total population size was only around 18 individuals. If the annual survival rate in the wild was 0.4 what is the probability that all 18 would have died in a single year?
Answer #1

Survivorship I = Ix bx

Where Ix is annual survival rate

bx is mortality rate

I = (0.4) (0.89) = 0.356

Probability of mortality f total population = 18/ 0.356 = 50.5%